# Graph-based Approaches and Functionalities in Retrieval-Augmented Generation: A Comprehensive Survey

### v1

@@ 4.2.1. A Unified View
In addition to utilizing existing knowledge graphs, it is also common to transform plain text into a knowledge graph through artificial means, which is known as open information extraction (OpenIE) (angeli2015leveraging, ; etzioni2008open, ; zhou2022survey, ). This process extracts key information—such as entities, relationships, and contextual meanings—from textual documents. An instruction-tuned language model then identifies these entities and their relationships to generate structured triples. As illustrated in Figure 3, given a paragraph describing the collaboration between Albert Einstein and Marcel Grossmann, the instruction-tuned LM captures the relationships among entities and converts them into structured triples. In longer contexts, multiple relationships emerge, linking entities to various neighbors and ultimately forming a complex graph.
@@ 4.2.2. Graph Generation Methods Based on Texts
KG-FiD (ref:kgfid, ) is a graph-based framework that constructs knowledge graphs directly from textual data to facilitate more effective information retrieval. Specifically, it begins by extracting text segments from the dataset and linking them to their associated entities. Using the BERT model for passage encoding, KG-FiD computes similarity scores and reranks the passages within entities to enhance retrieval accuracy. Additionally, a Graph Attention Network (GAT) is utilized to propagate information across the graph, leveraging relationships between entities to retrieve relevant node information. This approach ensures that the graph representation is both contextually rich and informative, leading to improved retrieval outcomes.
@@
MultiQG (ref:noname8, ) proposes a method to answer multi-hop complex queries over knowledge bases by generating query graphs with constraints. The approach integrates constraints early in the process to guide the construction of query graphs, effectively reducing the search space. Candidate query graphs are ranked based on their embedding similarity with the question using a neural network, ensuring accurate alignment. By combining beam search and semantic matching, the method attains state-of-the-art performance on standard benchmark datasets like ComplexWebQuestions, demonstrating significant improvements in precision and F1 scores. This framework addresses the limitations of traditional RAG systems, which struggle to handle constraints effectively, improving accuracy in complex question.
GLBK (ref:noname5, ) introduces a novel retrieval method for biomedical knowledge that leverages a knowledge graph to address the limitations of embedding-based similarity retrieval. This method constructs a biomedical knowledge graph by utilizing entity recognition and relationship extraction techniques, specifically fine-tuning PubmedBERT (ref:pubmedbert, ) for the biomedical domain. This graph builds text chunks from the PubMed  [^1] https://ftp.ncbi.nlm.nih.gov/pubmed/ corpus by associating them with nodes (entities like genes, diseases, or compounds) and edges (relationships). Unlike embedding-based retrieval methods, which tend to prioritize over-represented clusters in the data, this graph-based approach re-balances information by under-sampling dense clusters and ensuring access to long-tail, less frequent yet valuable knowledge. By combining the semantic strengths of embedding similarity with the structural insights from the knowledge graph, the proposed hybrid model demonstrates superior performance of accuracy, addressing the information overload problem prevalent in biomedical research retrieval.
ATLANTIC (ref:atlantic, ) is a structure-aware retrieval-augmented language model that addresses limitations in current RAG systems by integrating document structural relationships. A heterogeneous document graph is constructed from scientific literature, capturing four types of relationships: co-citation, co-topic, co-venue, and co-institution, to connect documents across 15 scientific disciplines. This structure is used to encode the relationships with a pre-trained Heterogeneous Graph Transformer (HGT) (hu2020heterogeneous, ), which generates structural embeddings. These embeddings are fused with semantic embeddings from text, enabling ATLANTIC to retrieve passages that are both semantically relevant and structurally coherent. The approach mitigates the shortcomings of text-only retrieval methods, particularly in interdisciplinary domains where relational context is crucial. The model demonstrates improved faithfulness and relevance in retrieving passages for scientific tasks like question answering and document classification.
GNN-Ret (ref:gnn-ret, ) constructs a database tailored for improved passage retrieval in complex question-answering scenarios. The authors create a ”Graph of Passages” (GoPs) by linking textual passages based on two key principles: structural relationships and shared keywords. Structural connections are established by linking adjacent passages in documents, maintaining their order and context. Keyword connections are derived by extracting entities using LLMs, and then linking passages that share the same keywords to highlight semantic associations. This graph-based representation addresses challenges like retrieving semantically distant but related passages, enabling improved retrieval coverage and more accurate question answering. The innovative use of GoPs mitigates the limitation of traditional database, which often treat passages as isolated units, and introduces an efficient way to model and retrieve interconnected information.
@@ 5. Graph-Driven R&P
Deterministic Algorithms Precise, reliable Intensive computation KG-GPT (ref:kg-gpt, ), GLBK (ref:noname5, ), GRAG (ref:grag, ), HyKGE (ref:hykge, ), Knowledge Solver (ref:ksl, ), KnowledGPT (ref:knowledgpt, ), GGE (ref:gge, ), Engine (ref:engine, ), GraphBridge (ref:graphbridge, ), MMGCN (ref:mmgcn, ), NLGraph (ref:nlgraph, ), GraphEval2000 (ref:GraphEval2000, )
Non-parameterized Algorithms Probabilistic Algorithms Scalable, strong adaptability Less precise, less interpretable HippoRAG (ref:hipporag, ), MindMap (ref:mindmap, ), OreoLM (ref:oreolm, ), MultiQG (ref:noname8, ), MINERVA (ref:minerva, ), MuseGraph (ref:musegraph, ), Walklm (ref:walklm, )
Heuristic-Based Algorithms Efficient, scalable Uncertain results depend on heuristics RA-SIM (ref:ra-sim, ), KG-Rank (ref:kg-rank, ), SKP (ref:skp, ), NuTrea (ref:nutrea, ), Zeshel (ref:zeshel, ), Conll (ref:conll, ), RNG-KBQA (ref:rng, ), ArcaneQA (ref:arcaneqa, ), HybridRAG (ref:hybridrag, ),MedGraphRAG (ref:medgraphrag, ), TOG2 (ref:tog2, ), DepsRAG (ref:depsrag, ), KELP (ref:kelp, ), KGQA (ref:kgqa, ), Graph-LLM (ref:Graph-LLM, ), GLBench (ref:GLBench, )
Learning-based Algorithms Convolutional-based Algorithms Efficient local structure modeling Limited global information GNN-RAG (ref:gnn-rag, ), MHKG (ref:feng, ), RA-SIM (ref:ra-sim, ), GenKGQA (ref:gentkgqa, ), GRAG (ref:grag, ), GNN-Ret (ref:gnn-ret, ), ETD (ref:etd, ), NuTrea (ref:nutrea, ), KAM-CoT (ref:kamcot, ), ConvE (ref:wn, ), REANO (ref:reano, ), SURGE (ref:SURGE, ), GraphGPT (ref:GraphGPT, )
Attention-Based Algorithms Global structure modeling High computational cost ATLANTIC (ref:atlantic, ), OpenSCR (ref:opencsr, ), GGE (ref:gge, ), GER (ref:ger, ), HSGE (ref:hsge, ), HamQA (ref:hamqa, ), KG-FiD (ref:kgfid, ), Fact (ref:fact, )
@@ 5.1.1. A Unified View
The non-parameterized methods in the retrieval process of RAG focus on extracting relevant knowledge from the graph using predefined rules, without involving any trainable parameters. Specifically, given the query $q$ and the knowledge graph $\mathcal{G}$, the retrieval response $R_{q}$ from the graph can be formulated as:
(1) $R_{q}=f(\mathcal{G},e(q)),$
where $e(\cdot)$ is the entity or relationship extraction function and $f(\cdot)$ is a deterministic function that follows predefined rules. To minimize overlaps of these deterministic functions and ensure the uniqueness of each group, we categorize the methods into the following three classes: Deterministic Graph Algorithms, Probabilistic Graph Algorithms, and Heuristic-Based Graph Algorithms. We outline the characteristics of these methods below and provide a visual representation in Figure 4.
$\bullet$ Deterministic Graph Algorithms. These algorithms focus on providing exact and reliable solutions to graph-related problems, ensuring correctness and precision. They prove especially effective for tasks requiring strict guarantees, such as subgraph isomorphism detection, shortest path computation, and graph traversals (e.g., Dijkstra’s algorithm (dijkstra2022note, ), Bellman-Ford algorithm (bellman1958routing, )). Their main advantage lies in their accuracy, but they may incur high computational costs for large-scale graphs. Figure 4 illustrates an example of the exact path computed between two nodes, highlighting how deterministic algorithms ensure optimality in path finding.
@@
Figure 4. Non-parameterized Algorithms for Graph Retrieval
@@ 5.1.2. Methods Using Deterministic Graph Algorithms
GLBK (ref:noname5, ) introduces a knowledge-graph-based retrieval method to address information overload in biomedical literature. To retrieve relevant text chunks for answering user queries, this method employs entity recognition to identify entities within the user question and constructs the shortest path linking these entities within the knowledge graph. The shortest path algorithm ensures the retrieval of text chunks associated with entities along the path and their neighboring edges, facilitating the generation of meaningful, non-trivial answers by the synthesizer LLM. This approach is especially effective in uncovering indirect relationships between entities, offering explanations and revealing potential discoveries. The use of graph distance metrics highlights the power of deterministic graph algorithms in rebalancing data and accessing rare but critical biomedical knowledge.
HyKGE (ref:hykge, ) introduces a novel hypothesis knowledge graph framework tailored to enhance medical LLM responses. The framework constructs a structured database by extracting entities and relations from user queries and hypothesis outputs generated by LLMs. Through named entity recognition (NER) and embedding alignment, the identified entities are connected to a medical knowledge graph. To solve the challenges of aligning unstructured queries with high-quality structured knowledge, HyKGE utilizes deterministic graph algorithms to retrieve three types of reasoning chains—paths, co-ancestor chains, and co-occurrence chains—between anchor entities. This enables precise and efficient retrieval of relevant, logical connections, addressing issues like noise and redundancy in retrieved data. The pruned reasoning chains are then integrated into LLMs, enhancing the accuracy and reliability of generated responses in complex medical scenarios.
GNN-RAG (ref:gnn-rag, ) combines the reasoning power of GNNs and the natural language processing abilities of LLMs for KGQA. The method begins with GNN reasoning over dense subgraphs to identify potential answer candidates. It then computes the shortest paths that connect question entities to the identified answers, which retrieve the exact information related to the candidate. These shortest paths represent meaningful reasoning paths within the knowledge graph, providing a transparent explanation for the reasoning process. The extracted paths are then verbalized into natural language and fed into an LLM to generate the final answer. This approach ensures accurate multi-hop reasoning while maintaining interpretability, significantly enhancing performance on complex KGQA tasks. Additionally, it leverages the efficiency of shortest-path algorithms to reduce computational complexity while ensuring that relevant reasoning paths are included.
KG-GPT (ref:kg-gpt, ) employs an entity matching in KGQA by segmenting sentences into sub-sentences aligned with triples. Specifically, using LLMs, it matches entities from the question to entities in the KG based on semantic similarity, creating precise entity pairs. This process identifies relations connecting these entities, forming a sub-KG. The sub-KG serves as evidence for reasoning tasks, enabling robust multi-hop inference. By integrating sentence segmentation, graph retrieval, and LLM inference, KG-GPT achieves high accuracy in KGQA, ensuring entity alignment and evidence precision for reliable reasoning and fact verification.
@@ 5.1.3. Methods Using Probabilistic Graph Algorithms
Walklm (ref:walklm, ) serves as a standardized fine-tuning framework for language models tailored to attributed graph embedding, providing a new approach for integrating the LLMs with graph retrieval tasks. The core of WalkLM involves a probabilistic method for graph retrieval that leverages random walks. Specifically, the framework starts by sampling random walks from a given graph, where each walk begins from a selected node and continues by traversing edges randomly. The nodes and edges encountered in the random walks are then converted into text, effectively transforming the graph structure into sequences that can be processed by LLMs. This text is further tokenized to create a list of tokens for embedding formulation. By fine-tuning a pre-trained language model on these graph-derived textual sequences, WalkLM generates embeddings that encapsulate both the attribute semantics and structural relationships inherent in the graph. This method effectively unifies structural and attribute-based information into a meaningful representation that can be used for a variety of downstream tasks, such as node classification and link prediction.
HippoRAG (ref:hipporag, ) extracts entities from a query using LLMs and maps them to nodes in a knowledge graph based on cosine similarity. It then applies the Personalized PageRank (PPR) algorithm, seeded with these query nodes, to distribute probabilities across the graph and identify relevant neighborhoods. To refine retrieval accuracy, the query node probabilities are adjusted using a node specificity metric before PPR, enhancing the ranking of indexed passages and enabling precise multi-hop reasoning in a single step.
@@ 5.1.4. Methods Using Heuristic-Based Graph Algorithms
HybridRAG (ref:hybridrag, ) integrates vector and graph retrieval methods to improve the analysis and use of financial documents. By utilizing the strengths of both approaches, HybridRAG provides a more comprehensive retrieval process. For vector-based knowledge, it utilizes the LangChain framework  [^2] https://www.langchain.com/ to query a vector store and retrieve the document segments most pertinent to a given question. On the other hand, for graph-based retrieval, it employs the GraphQAChain from LangChain to traverse the knowledge graph, specifically accessing 1-hop neighbors of entities detected in the question. Compared to traditional vector-based retrieval alone, this heuristic graph algorithm is straightforward and contributes to a deeper context understanding, thereby improving the quality of generated answers.
GRAG (ref:grag, ) uses heuristic algorithms to retrieve $K$-hop ego-graphs as candidate subgraphs for graph retrieval. It identifies key ego-graphs based on cosine similarity between query embeddings and ego-graph embeddings, which are derived through mean pooling of textual embeddings generated by a pre-trained SentenceBERT (ref:sentence-bert, ) model. To refine the selection, a learnable pruner adaptively masks irrelevant nodes and edges using MLPs, which assign relevance scores based on proximity to the query. The top-ranked subgraphs are then merged into an optimal subgraph structure, enabling efficient retrieval while preserving topological and semantic relevance for downstream reasoning tasks.
@@ 5.2.1. A Unified View
In contrast to non-parameterized approaches, the learning-based retrieval method can be formulated as a parameterized function $f_{\theta}(\cdot)$ that learns to retrieve relevant knowledge by optimizing an objective function:
(2) $R_{q}=f_{\theta}(\mathcal{G},e(q)),$
where $\theta$ represents the learnable parameters within $f_{\theta}(\cdot)$. These parameters are trained on a dataset using an optimization objective, enabling the retrieval mechanism to adapt dynamically to different data distributions and retrieval tasks. Unlike traditional rule-based retrieval, which relies on predefined heuristics, learning-based approaches can generalize across domains, capture complex semantic relationships, and improve retrieval relevance over time. Depending on how the system processes and extracts information from the knowledge graph, learning-based retrieval algorithms can be categorized into the following two major classes:
$\bullet$ Convolutional-based Algorithms. Convolutional-based algorithms are designed to exploit the structural properties of graphs by performing neighborhood aggregation through trainable network layers. These methods are inspired by convolutional operations in image processing but are adapted to graphs to capture local connectivity patterns and facilitate information propagation across nodes. A representative example is the Graph Convolutional Network (GCN) (ref:gcn, ), which employs a layer-wise message-passing mechanism to iteratively aggregate features from neighboring nodes. As shown on the LHS of Figure 5, given the central node $x_{u}$ and its neighbors $x_{1}\sim x_{3}$, different weights $C_{uu}$ and $C_{u1}\sim C_{u3}$ are applied to aggregate the features from each neighbor. This weighted aggregation allows the central node to incorporate diverse contextual information, leading to richer feature representations. By stacking multiple layers, these algorithms allow retrieval models to learn hierarchical representations, enhancing their ability to extract relevant knowledge from graph-structured data.
$\bullet$ Attention-Based Algorithms. Attention-based algorithms in graph retrieval have gained significant interest due to their ability to dynamically prioritize the most relevant parts of the graph during learning. Unlike traditional approaches that treat all neighboring nodes equally, these methods employ attention mechanisms to assign varying importance weights ($A_{uu}$ and $A_{u1}\sim A_{u3}$ in RHS of Figure 5) to nodes and edges, enabling the model to focus on key relationships that are most critical for a given task. A prominent example is the Graph Attention Network (GAT) (ref:gat, ), which introduces node-wise attention coefficients to aggregate information more selectively. This adaptive weighting mechanism allows attention-based retrieval models to capture long-range dependencies, refine contextual understanding, and improve retrieval precision.
Figure 5. Learning-based Algorithms

### v2

@@ 4.2.1. A Unified View
In addition to reusing existing knowledge graphs, we consider a broader class of graph-powered databases constructed from text. A common route is transforming plain text into a knowledge graph through artificial means, which is known as open information extraction (OpenIE) (Angeli et al., 2015; Etzioni et al., 2008; Zhou et al., 2022). This process extracts key information—such as entities, relationships, and contextual meanings—from textual documents. An instruction-tuned language model then identifies these entities and their relationships to generate structured triples. As illustrated in Figure 4, given a paragraph describing the collaboration between Albert Einstein and Marcel Grossmann, the instruction-tuned LM captures the relationships among entities and converts them into structured triples. In longer contexts, multiple relationships emerge, linking entities to various neighbors and ultimately forming a complex graph.
Beyond vanilla knowledge graphs, we also include multiple forms graphs distilled from text: (1) document-level structures such as entity–sentence and sentence–sentence links (HippoRAG (Gutiérrez et al., 2024), ATLANTIC (Munikoti et al., 2023), GNN-Ret (Li et al., 2024f), OpenSCR (Han et al., 2023), FiD (Yu et al., 2022), MedGraphRAG (Wu et al., 2024b)); (2) hierarchical graphs with each node as a subgraph (DALK (Li et al., 2024a), GER (Wu et al., 2023a)); (3) rule-based graph where each node aggregates multiple entities or facts consistent with a defined rule schema (FABULA (Ranade and Joshi, 2023), StarRAG (Zhu et al., 2025), ReTraCk (Chen et al., 2021)). These constructions produce property graphs with additional information and traceable evidence, enabling unique feature extractions from the knowledge base.
@@ 4.2.2. Graph Generation Methods Based on Texts
KG-FiD (Yu et al., 2022) is a graph-based framework that constructs knowledge graphs directly from textual data to facilitate more effective information retrieval. Specifically, it begins by extracting text segments from the dataset and linking them to their associated entities. Using the BERT model for passage encoding, KG-FiD computes similarity scores and reranks the passages within entities to enhance retrieval accuracy. Additionally, a Graph Attention Network (GAT) is utilized to propagate information across the graph, leveraging relationships between entities to retrieve relevant node information. This approach ensures that the graph representation is both contextually rich and informative, leading to improved retrieval outcomes.
@@
MultiQG (Lan and Jiang, 2020) proposes a method to answer multi-hop complex queries over knowledge bases by generating query graphs with constraints. The approach integrates constraints early in the process to guide the construction of query graphs, effectively reducing the search space. Candidate query graphs are ranked based on their embedding similarity with the question using a neural network, ensuring accurate alignment. By combining beam search and semantic matching, the method attains state-of-the-art performance on standard benchmark datasets like ComplexWebQuestions, demonstrating significant improvements in precision and F1 scores. This framework addresses the limitations of traditional RAG systems, which struggle to handle constraints effectively, improving accuracy in complex question.
@@ 5. Graph-Driven R&P
Deterministic Algorithms Precise, reliable Intensive computation KG-GPT (Kim et al., 2023a), GLBK (Delile et al., 2024), GRAG (Hu et al., 2024), HyKGE (Jiang et al., 2024c), Knowledge Solver (Feng et al., 2023), KnowledGPT (Wang et al., 2023c), GGE (Gao et al., 2022), Engine (Zhu et al., 2024c), GraphBridge (Wang et al., 2024c), MMGCN (Wei et al., 2019), NLGraph (Wang et al., 2024a), GraphEval2000 (Wu et al., 2024c)
Non-parameterized Algorithms Probabilistic Algorithms Scalable, strong adaptability Less precise, less interpretable HippoRAG (Gutiérrez et al., 2024), MindMap (Wen et al., 2024), OreoLM (Hu et al., 2022), MultiQG (Lan and Jiang, 2020), MINERVA (Das et al., 2018), MuseGraph (Tan et al., 2024a), Walklm (Tan et al., 2024b)
Heuristic-Based Algorithms Efficient, scalable Uncertain results depend on heuristics RA-SIM (Peng and Yang, 2024), KG-Rank (Yang et al., 2024a), SKP (Dong et al., 2023a), NuTrea (Choi et al., 2023), Zeshel (Logeswaran et al., 2019), Conll (Hoffart et al., 2011), RNG-KBQA (Ye et al., 2021), ArcaneQA (Gu and Su, 2022), HybridRAG (Sarmah et al., 2024),MedGraphRAG (Wu et al., 2024b), TOG2 (Ma et al., 2024), DepsRAG (Alhanahnah et al., 2024), KELP (Liu et al., 2024b), KGQA (Sen et al., 2023), Graph-LLM (Chen et al., 2024c), GLBench (Li et al., 2024d)
Learning-based Algorithms Convolutional-based Algorithms Efficient local structure modeling Limited global information GNN-RAG (Mavromatis and Karypis, 2024), MHKG (Feng et al., 2020), RA-SIM (Peng and Yang, 2024), GenKGQA (Gao et al., 2024b), GRAG (Hu et al., 2024), GNN-Ret (Li et al., 2024f), ETD (Liu et al., 2024a), NuTrea (Choi et al., 2023), KAM-CoT (Mondal et al., 2024), ConvE (Dettmers et al., 2018), REANO (Fang et al., 2024), SURGE (Kang et al., 2022), GraphGPT (Tang et al., 2024)
Attention-Based Algorithms Global structure modeling High computational cost ATLANTIC (Munikoti et al., 2023), OpenSCR (Han et al., 2023), GGE (Gao et al., 2022), GER (Wu et al., 2023a), HSGE (Sun et al., 2023), HamQA (Dong et al., 2023b), KG-FiD (Yu et al., 2022), Fact (Opsahl, 2024)
@@ 5.1.1. A Unified View
The non-parameterized methods in the retrieval process of RAG focus on extracting relevant knowledge from the graph using predefined rules, without involving any trainable parameters. Specifically, given the query $q$ and the knowledge graph $\mathcal{G}$, the retrieval response $R_{q}$ from the graph can be formulated as: $R_{q}=f(\mathcal{G},e(q)),$ where $e(\cdot)$ is the entity or relationship extraction function and $f(\cdot)$ is a deterministic function that follows predefined rules. To minimize overlaps of these deterministic functions and ensure the uniqueness of each group, we categorize the methods into the following three classes: Deterministic Graph Algorithms, Probabilistic Graph Algorithms, and Heuristic-Based Graph Algorithms. We outline the characteristics of these methods below and provide a visual representation in Figure 5.
$\bullet$ Deterministic Graph Algorithms. These algorithms focus on providing exact and reliable solutions to graph-related problems, ensuring correctness and precision. They prove especially effective for tasks requiring strict guarantees, such as subgraph isomorphism detection, shortest path computation, and graph traversals (e.g., Dijkstra’s algorithm (Dijkstra, 2022), Bellman-Ford algorithm (Bellman, 1958)). Their main advantage lies in their accuracy, but they may incur high computational costs for large-scale graphs. Figure 5 illustrates an example of the exact path computed between two nodes, highlighting how deterministic algorithms ensure optimality in path finding.
@@
Figure 5. Non-parameterized Algorithms and learning-based for Graph Retrieval
@@ 5.1.2. Methods Using Deterministic Graph Algorithms
GLBK (Delile et al., 2024) recognizes entities in the question, computes shortest paths that connect them in the knowledge graph, and retrieves text linked to the nodes and incident edges along these paths, which reveals indirect relations that support answer synthesis. HyKGE (Jiang et al., 2024c) aligns query and hypothesis entities to a medical graph and then collects three deterministic reasoning chains between anchor entities, namely simple paths, co-ancestor chains, and co-occurrence chains, followed by pruning that keeps the most relevant connections as retrieval evidence. GNN-RAG (Mavromatis and Karypis, 2024) first scores candidate answers with graph neural reasoning and then extracts the shortest paths from the question entities to these candidates so that the retrieved paths provide faithful and efficient multi hop evidence. KG-GPT (Kim et al., 2023a) segments the question into triple level units, matches entities to the graph through semantic alignment, identifies the relations that connect them, and assembles a subgraph that serves as compact retrieval support for downstream reasoning.
@@ 5.1.3. Methods Using Probabilistic Graph Algorithms
Within the taxonomy in Table 2, probabilistic graph algorithms cast retrieval as stochastic evidence flow on the graph. Walklm (Tan et al., 2024b) introduces targeted random walk sampling from query anchored nodes, which turns the visited trails into textual sequences and aggregates path evidence to rank candidate neighborhoods and passages. HippoRAG (Gutiérrez et al., 2024) illustrates diffusion-driven retrieval, where detected query entities are aligned to knowledge graph nodes and used as seeds for Personalized PageRank, which spreads probability through nearby structure and then applies specificity-based reweighting to select high probability neighborhoods and surface the associated passages for multi-hop reasoning.
@@ 5.1.4. Methods Using Heuristic-Based Graph Algorithms
HybridRAG (Sarmah et al., 2024) pairs vector search over the document store with a one-hop neighborhood walk from entities detected in the question, which enriches the retrieved context and improves answer quality while keeping the search path simple. GRAG (Hu et al., 2024) illustrates structured candidate selection, where K-hop ego graphs are ranked by embedding similarity to the query. Then a learnable pruner suppresses low-value nodes and edges, and the top candidates are merged into a compact subgraph for generation, which preserves semantic focus and graph topology under tight runtime budgets.
@@ 5.2.1. A Unified View
In contrast to non-parameterized approaches, the learning-based retrieval method can be formulated as a parameterized function $f_{\theta}(\cdot)$ that learns to retrieve relevant knowledge by optimizing an objective function: $R_{q}=f_{\theta}(\mathcal{G},e(q)),$ where $\theta$ represents the learnable parameters within $f_{\theta}(\cdot)$. These parameters are trained on a dataset using an optimization objective, enabling the retrieval mechanism to adapt dynamically to different data distributions and retrieval tasks. Unlike traditional rule-based retrieval, which relies on predefined heuristics, learning-based approaches can generalize across domains, capture complex semantic relationships, and improve retrieval relevance over time. Depending on how the system processes and extracts information from the knowledge graph, learning-based retrieval algorithms can be categorized into the following two major classes:
$\bullet$ Convolutional-based Algorithms. Convolutional-based algorithms are designed to exploit the structural properties of graphs by performing neighborhood aggregation through trainable network layers. These methods are inspired by convolutional operations in image processing but are adapted to graphs to capture local connectivity patterns and facilitate information propagation across nodes. A representative example is the Graph Convolutional Network (GCN) (Kipf and Welling, 2017), which employs a layer-wise message-passing mechanism to iteratively aggregate features from neighboring nodes. As shown on the LHS of Figure 5, given the central node $x_{u}$ and its neighbors $x_{1}\sim x_{3}$, different weights $C_{uu}$ and $C_{u1}\sim C_{u3}$ are applied to aggregate the features from each neighbor. This weighted aggregation allows the central node to incorporate diverse contextual information, leading to richer feature representations. By stacking multiple layers, these algorithms allow retrieval models to learn hierarchical representations, enhancing their ability to extract relevant knowledge from graph-structured data.
$\bullet$ Attention-Based Algorithms. Attention-based algorithms in graph retrieval have gained significant interest due to their ability to dynamically prioritize the most relevant parts of the graph during learning. Unlike traditional approaches that treat all neighboring nodes equally, these methods employ attention mechanisms to assign varying importance weights ($A_{uu}$ and $A_{u1}\sim A_{u3}$ in RHS of Figure 5) to nodes and edges, enabling the model to focus on key relationships that are most critical for a given task. A prominent example is the Graph Attention Network (GAT) (Veličković et al., 2018), which introduces node-wise attention coefficients to aggregate information more selectively. This adaptive weighting mechanism allows attention-based retrieval models to capture long-range dependencies, refine contextual understanding, and improve retrieval precision.